# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 3 Structure functions and dissipation
Figure 1: Dependence of the velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ (left), vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ (centre) and helicity structure function $\left\langle{\delta h}\right\rangle$ (right) on the Deborah number. For $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ the lines for the different $De$ are vertically shifted for increase the clarity.
@@
The velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ clearly shows the multiscaling behaviour of polymeric flows (Zhang et al., 2021; Rosti et al., 2023; Singh & Rosti, 2024). In the Newtonian case, and for $De\leq 1/3$, we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$ in the intermediate range of scales in agreement with the Kolmogorov predictions (Kolmogorov, 1941). For small $De$, i.e. in the limit of $\tau_{p}\rightarrow 0$, the polymers do not stretch, and they only marginally influence the velocity structure function. For $De\approx 1$ the polymeric relaxation time is comparable with the large time scale of the flow and the polymers interact with the energy cascade, inducing a steeper depletion of $\left\langle{\delta q^{2}}\right\rangle$ with $r$ in the so-called elastic range of scales (see for example De Angelis et al., 2005; Perlekar et al., 2010). For $De=1$ we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ that matches the energy spectrum scaling of $\kappa^{-2.3}$ (Rosti et al., 2023), where $\kappa$ is the wavenumber. Due to large relaxation times, the polymers now remain stretched when further increasing $De$ and begin to decouple from the carrier flow. For $De\geq 3$, therefore, the influence of the polymers weakens, and $\left\langle{\delta q^{2}}\right\rangle$ shows a less steeper dependence on $r$, progressively recovering the $r^{2/3}$ Kolmogorov scaling. For $De=9$ we measure $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.1}$ over a relatively wide range of scales. Note that, $\left\langle{\delta q^{2}}\right\rangle\sim r^{2}$ at the small scales $r\rightarrow 0$ for all cases, which is known to follow from the smoothness of the velocity field at small scales (Schumacher et al., 2007). With the real space structure function, indeed, we can not observe the increase of the energy in the deep dissipative range that is found in the Fourier space energy spectrum by Perlekar et al. (2010), Rosti et al. (2023), and by Singh & Rosti (2024) with the present database.
@@
The vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ is shown in the central panel of figure 1. It is clear that enstrophy growth (generation) is confined to small scales for all $De$, as already observed by other authors for HIT (see for example Jiménez et al., 1993; Ishihara et al., 2013; Elsinga et al., 2017). Indeed, $\left\langle{\delta\omega^{2}}\right\rangle$ rapidly increases with $r$ at the small scales and saturates in the inertial range. As detailed in the following, the polymers reduce the vorticity fluctuations at the small scales (Liberzon et al., 2005, 2006; Perlekar et al., 2010) which results in a decrease of $\left\langle{\delta\omega^{2}}\right\rangle$ at all $r$. Note that, like for $\left\langle{\delta q^{2}}\right\rangle$, the effect of the polymers on the vorticity structure function is non-monotonous, with $\left\langle{\delta\omega^{2}}\right\rangle$ being minimum for $De=3$ at all scales.
@@
The dependence of the helicity structure functions $\left\langle{\delta h}\right\rangle$ on $De$ is shown in the right panel of figure 1. In HIT, we find $\left\langle{\delta h}\right\rangle\sim r^{2/3}$ in the inertial range, in agreement with the existence of the cascade of helicity from larger to the smaller scales envisaged by Brissaud et al. (1973); see also Polifke & Shtilman (1989), Borue & Orszag (1997) and Baj et al. (2022) and section §6. Note also that $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta q^{2}}\right\rangle$ have an almost identical $r^{2/3}$ range, indicating that the inertial range for helicity has an almost identical span across scales as that for energy, confirming the findings of Chen et al. (2003a). The effect of the polymers on $\left\langle{\delta h}\right\rangle$ resembles what has been observed for $\left\langle{\delta q^{2}}\right\rangle$, suggesting that the two cascades are influenced in a similar fashion. For small and large $De$, the influence of the polymers is rather small: the $r^{2/3}$ scaling is recovered for $De\rightarrow 0$ and $De\rightarrow\infty$. For intermediate $De$, instead, $\left\langle{\delta h}\right\rangle$ shows a steeper dependence on $r$ in the elastic range of scales. For $De=1$, we find $\left\langle{\delta h}\right\rangle\sim r^{0.9}$, which is shallower compared to the $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ scaling for the energy, hinting a weaker influence of polymers on the helicity cascade. Overall, the results show that helicity is mainly concentrated at the large scales for all $De$: turbulence cascade tends to restore the mirror symmetry at the small scales in both Newtonian and polymeric turbulence (Ditlevsen & Giuliani, 2001; Chen et al., 2003a; Baj et al., 2022); see §6 for additional discussion.
To provide further insights on the scale-by-scale influence of the polymers on the energy and enstrophy distributions, we consider the signature function $V(r)$, introduced by Davidson (2004) and Davidson & Pearson (2005) to eliminate the enstrophy contribution of eddies of scales larger than $r$ from the second-order structure function, i.e.
$V(r)=-\frac{1}{2}r^{2}\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial}{\partial r}\left(\frac{3}{4}S_{2}(r)\right)\right).$
The signature function has the following properties: (i) $\int_{0}^{r}V(r)\text{d}r\geq 0$ and (ii) $\int_{0}^{\infty}V(r)\text{d}r=\frac{1}{2}\left\langle{\bm{u}^{2}}\right\rangle$(Davidson, 2004). Therefore, $V(r)$ is interpreted as the energy density and $rV(r)$ estimates the kinetic energy associated with eddies of scale $r$. Similarly, $10V(r)/r$ is a measure of the enstrophy associated with eddies of size $r$.
Figure 2: Contribution of eddies of size $r$ on the (left) total energy and on the (centre) total enstrophy for different Deborah numbers. $V(r)$ is the signature function introduced by Davidson (2004) and Davidson & Pearson (2005). The right panel shows the distribution of the helicity $h=\bm{u}\cdot\bm{\omega}$ for different $De$.
The left and middle panels of figure 2 show the scale-by-scale distributions of $rV(r)$ and $10V(r)/r$ for different $De$. We start with the energy distribution in the left panel. As expected, largest scales are the most energetic, and $rV(r)$ diminishes as $r$ decreases, being null for $r\rightarrow 0$. The influence of the polymers is felt more at the small scales, while the amount of energy associated with the largest scales remains largely unaffected. This is consistent with the observations in earlier studies (Bhattacharjee & Thirumalai, 1991; Fouxon & Lebedev, 2003; De Angelis et al., 2005; Perlekar et al., 2010) that polymers decrease the energy content at small and intermediate scales, with $r_{p}$, that separates the large and intermediate scales, changing with $De$. In agreement with the findings of Rosti et al. (2023), the energy depletion at the small scales does not show a monotonous dependence on $De$. Interestingly, the effect of $De$ on $rV(r)$ changes with $r$, suggesting that the local coupling between the fluid and the polymers depends on the local ratio between the polymeric relaxation time and the characteristic time scale of the flow at that scale, i.e. $\tau_{p}/\tau_{f}(r)$, where $\tau_{f}(r)$ decreases at smaller scales (for example, K41 implies that the flow time scales in Newtonian turbulence decay as $\tau_{f}(r)\sim r^{2/3}$). To understand this better, let us focus on $1\leq De\leq 9$. At $r\approx 0.02$ the energy depletion is maximum for $De=3$ and minimum for $De=9$, while for $r\approx 0.6$ it is maximum for $De=9$ and minimum for $De=1$: the inversion of the trend occurs first at the smaller scales, where indeed the local $\tau_{p}/\tau_{f}(r)$ is larger and the decoupling between the fluid and the polymers occurs for smaller values of $\tau_{p}$. A last comment regards the overshoot of $rV(r)$, i.e. the gentle local increase of energy observed for scales between the inertial and the far-dissipation ranges, usually referred to as bottleneck (see for example Falkovich, 1994; Lohse & Müller-Groeling, 1995; Frisch et al., 2013). Although a detailed investigation of this effect is not within the scope of the present work, it is worth noticing that when the coupling between the fluid and the polymers is maximum, i.e. $De\approx 1$, this overshoot is annihilated and the local increase of $rV(r)$ is not observed.
We now briefly move to the distribution of $10V(r)/r$, shown in the central panel of figure 2. It confirms that the enstrophy is limited to the small scales for all $De$, and that the influence of the polymers in negligible at the largest scales. Akin to energy, and in agreement with the distribution of $\left\langle{\delta\omega^{2}}\right\rangle$, we observe that the influence of the polymers on the scale-by-scale enstrophy is not monotonic in $De$: the maximum enstrophy reduction is observed for $De=3$ at the small scales and for $De=9$ at larger scales.
The right panel of figure 2 describes the influence of the polymers on the probability distribution function (pdf) of $h=\bm{u}\cdot\bm{\omega}$. Helicity is not positive definite, and has an asymmetric distribution for all cases; it is right skewed and the right tail is longer for all $De$. In agreement with the distribution of $\left\langle{\delta h}\right\rangle$ shown in figure 1, the presence of the polymers narrows the distribution of $h$, suggesting that polymers reduce the extreme events of helicity. However, the asymmetry of the distribution increases, and even more so as $De$ increases, indicating that the presence of the polymers favours events that break mirror symmetry. Again, the $De$ dependence is not monotonous, with an inversion of the effects of the polymers on the tails observed for $De\geq 3$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
We show the nature of dissipation of all three quantities in figure 3. These quantities play a key role in the scale-by-scale cascades. In Newtonian turbulence, for the two averaged, inviscid invariants (energy and helicity) the dissipation rate matches the injection rate at the large scales and the averaged total flux in the inertial range of scales. The distributions of the dissipations show large tails, which are indicative of small-scale intermittency (Sreenivasan & Antonia, 1997). As expected, the distributions narrow for the polymeric case: polymers reduce the magnitude of the velocity derivatives, thus largely reducing the regions of large dissipations (Liberzon et al., 2005, 2006). The distributions of the energy and enstrophy dissipation collapse nicely for all $De$ once the quantities are normalised with the respective standard deviation, similarly to what was shown by Perlekar et al. (2010) and is consistent with the $De$ invariance of intermittency corrections observed in Rosti et al. (2023).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 4: Distribution of the the polymer/fluid exchange terms $\pi$ for $\delta q^{2}$, $\delta\omega^{2}$ and $\delta h$ for different Deborah numbers. Left: $\pi^{\delta q^{2}}$, centre: $\pi^{\delta\omega^{2}}$ and right: $\pi^{\delta h}$. The bottom panels plot the distribution of the three quantities normalised with their root-mean-square value.
@@
We conclude this section with the distributions of $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$ and $\pi^{\delta h}$, which capture the additional sink/source of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ due to the polymeric microstructure. These quantities are not positive definite, meaning that locally polymers may act as either a sink or a source for the velocity/vorticity fluctuations. The distribution of $\pi^{\delta q^{2}}$ is right-skewed for all $De$, with the highly right skewed distribution confirming the positive average value $\langle\pi^{\delta q^{2}}\rangle=\langle\epsilon_{p}\rangle$ (De Angelis et al., 2005). As $De$ increases, both the right and left tails widen, meaning that sink and source intense events are equally promoted by the larger stretching of the polymers. The effect of $De$ is much more significant on $\pi^{\delta\omega^{2}}$, whose distributions are shown in the central panels of figure 4. They become increasingly left skewed with increasing $De$. At small $De<1$, the distributions are right skewed and, on average, the polymers act as a sink for the enstrophy. For $De\geq 1$, instead, the distribution of $\pi^{\delta\omega^{2}}$ is left skewed. In this case, due to the larger stretching of the polymers, the fluid-polymer interaction facilitates intense production of enstrophy. As detailed in §5, this qualitative change is rooted in a profound re-organisation of the small-scale fluctuations. Eventually, the right panels in figure 4 shows that the distribution of $\pi^{\delta h}$ is symmetric, and that both tails become longer as $De$ increases.
4 Scale energy

### v2

@@ 3 Structure functions and dissipation
Figure 1: Instantaneous visualisations of energy, enstrophy and helicity fields for $De=0$ (top) and $De=1$ (bottom). Left: $q^{2}=u_{i}u_{i}$. Centre: $\omega^{2}=\omega_{i}\omega_{i}$. Right: $h=u_{i}\omega_{i}$. All quantities are normalised with the average value in the considered slice.
Figure 1 presents representative two-dimensional slices of the three key quantities: kinetic energy $q^{2}=u_{i}u_{i}$ (left), enstrophy $\omega^{2}=\omega_{i}\omega_{i}$ (centre) and helicity $h=u_{i}\omega_{i}$ (right), obtained from both Newtonian $De=0$ (top) and visoelastic $De=1$ (bottom) simulations. Enstrophy and helicity exhibit smaller-scale, more intermittent structures due to their direct dependence on velocity gradients. The addition of polymers significantly modifies the flow structures. In the viscoelastic case, all three fields show noticeably smoother small-scale features, reflecting the polymer-induced suppression of small-scale fluctuations.
@@
Refer to captionRefer to caption
Figure 2: Dependence of the velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ (left), vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ (centre) and helicity structure function $\left\langle{\delta h}\right\rangle$ (right) on the Deborah number. For $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ the lines for the different DeDe are vertically shifted for increase the clarity.
@@
The velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ clearly shows the multiscaling behaviour of polymeric flows (Zhang et al., 2021; Rosti et al., 2023; Singh & Rosti, 2025). In the Newtonian case and for $De\leq 1/3$ we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$ in the intermediate range of scales in agreement with the Kolmogorov predictions (Kolmogorov, 1941). For small $De$, i.e. in the limit of $\tau_{p}\rightarrow 0$, the polymers do not stretch, and they only marginally influence the velocity structure function. For $De=\mathcal{O}(1)$ the polymeric relaxation time is comparable with the large time scale of the flow and the polymers interact with the energy cascade, inducing a steeper depletion of $\left\langle{\delta q^{2}}\right\rangle$ with $r$ in the so-called elastic range of scales (De Angelis et al., 2005; Perlekar et al., 2010). For $De\geq 1$ we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ that matches the energy spectrum scaling $E(k)\sim k^{-2.3}$ (Rosti et al., 2023), $k$ being the wavenumber. At large Deborah numbers ($De\gg 1$), polymers possess long relaxation times and tend to remain stretched, progressively decoupling from the carrier flow. As a result, their influence diminishes and the scaling of $\left\langle{\delta q^{2}}\right\rangle$ reverts toward the classical Kolmogorov prediction $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$ as $De\rightarrow\infty$ (Rosti et al., 2023). Note that Kolmogorov scaling is not fully recovered for the largest $De$ considered in the current parameter set (see §A).
The scenario is thus the following. At low Deborah numbers $De\ll 1$, the polymers behave essentially as stiff chains and undergo only limited stretching. Consequently, the fluctuations in the polymer conformation tensor $R_{ij}$ and the polymer stress tensor $T_{ij}$ are rather small, resulting in negligible non-Newtonian stresses $\partial T_{ij}/\partial x_{j}\approx 0$. Under these conditions, the flow statistics closely resemble those of a purely Newtonian fluid, with the energy spectrum and structure functions exhibiting the classical $E(k)\sim k^{-5/3}$ and $S_{p}(r)\sim r^{p/3}$ scalings, where $S_{p}\equiv\left\langle{\left(\left(\bm{u}(\bm{x}+\bm{r},t)-\bm{u}(\bm{x},t)\right)\cdot\bm{r}/r\right)^{p}}\right\rangle$, with $r=|\bm{r}|$; in HIT $S_{2}\sim\left\langle{\delta q^{2}}\right\rangle$ due to the incompressibility constraint (Hill, 1997). As $De$ increases towards $\mathcal{O}(1)$, the scaling behaviour changes. In this intermediate regime, the polymer relaxation time becomes comparable to the characteristic time scale of the fluid. Here, the energy transfer is mediated by the polymer elasticity rather than solely by the nonlinear cascade (see §4.1), leading to the modified scalings $E(k)\sim k^{-2.3}$ for the energy spectrum and $S_{2}(r)\sim r^{1.3}$ for the second-order structure function. However, this regime cannot persist in the limit $De\rightarrow\infty$. For $De\gg 1$, indeed, the polymers effectively decouple from the fluid dynamics. Due to their long relaxation times, the polymers respond to flow changes with significant delay, and remain stretched for extended periods. This suppresses fluctuations in polymer lengths, weakens the non-Newtonian stresses $\partial T_{ij}/\partial x_{j}$, and causes the flow to revert towards a Kolmogorov-like behaviour. The progressive narrowing of the scale range associated with the elastic scaling for $De\gtrsim 1$ is more clearly observed in Fourier space, as discussed in appendix §A.
The exact value of $De$ at which the flow transitions between different regimes is influenced by the method used to sustain turbulence. For instance, Rosti et al. (2023) conducted simulations at the same Reynolds number as in the present study but adopted the forcing scheme of Eswaran & Pope (1988)—which injects energy randomly at low wavenumbers—, rather than the ABC forcing used here—which generates a more coherent large-scale flow structure (Chiarini et al., 2024). With their approach, Rosti et al. (2023) observed the onset of Kolmogorov-like scaling already at $De=9$, a behaviour not seen in our dataset. Distinct forcing mechanisms adopted to sustain turbulence indeed affect differently the large-scale dynamics and, therefore, the characteristic time scale used to define $De$. As a result, the precise values and range of $De$ over which polymers dominate, and the $S_{2}\sim r^{1.3}$ scaling emerges, varies slightly depending on the external forcing. Nevertheless, the overall qualitative behaviour remains unchanged, with only modest $\mathcal{O}(1)$ shifts in the critical $De$ values delineating the different regimes.
It is worth noting that $\left\langle{\delta q^{2}}\right\rangle\sim r^{2}$ as $r\rightarrow 0$ across all cases, consistent with the expected smoothness of the velocity field at small scales (Schumacher et al., 2007). In real space, this smooth behaviour precludes the observation of the rise in energy at deep dissipative scales that has been reported in the Fourier-space energy spectra by Perlekar et al. (2010), Rosti et al. (2023), and by Singh & Rosti (2025) using the present dataset.
@@
The vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ is shown in the central panel of figure 2. It is clear that enstrophy growth (generation) is confined to small scales for all $De$, as already observed by other authors for Newtonian HIT (Jiménez et al., 1993; Ishihara et al., 2013; Elsinga et al., 2017). Indeed, $\left\langle{\delta\omega^{2}}\right\rangle$ rapidly increases with $r$ at the small scales and saturates in the inertial range. The polymers reduce the vorticity fluctuations at the small scales (Liberzon et al., 2005, 2006; Perlekar et al., 2010) which results in a decrease of $\left\langle{\delta\omega^{2}}\right\rangle$ at all $r$. Note that, like for $\left\langle{\delta q^{2}}\right\rangle$, the effect of the polymers on the vorticity structure function is non-monotonous, with $\left\langle{\delta\omega^{2}}\right\rangle$ being minimum for $De=3$ at all scales.
@@
The dependence of the helicity structure functions $\left\langle{\delta h}\right\rangle$ on $De$ is shown in the right panel of figure 2. In the Newtonian case, we find $\left\langle{\delta h}\right\rangle\sim r^{2/3}$ in the inertial range, in agreement with the existence of the cascade of helicity from larger to the smaller scales envisaged by Brissaud et al. (1973); see also Polifke & Shtilman (1989), Borue & Orszag (1997) and Baj et al. (2022) and §4.3. Note also that $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta q^{2}}\right\rangle$ have an almost identical $r^{2/3}$ range, indicating that the inertial range for helicity spans nearly the same range of scales as that for energy, confirming the findings of Chen et al. (2003a). The effect of the polymers on $\left\langle{\delta h}\right\rangle$ resembles what has been observed for $\left\langle{\delta q^{2}}\right\rangle$, suggesting that the two cascades are influenced in a similar fashion. For small and large $De$, the influence of the polymers is rather small: the $r^{2/3}$ scaling is recovered for $De\rightarrow 0$ and $De\rightarrow\infty$. For intermediate $De$, instead, $\left\langle{\delta h}\right\rangle$ shows a steeper dependence on $r$ in the elastic range of scales. For $De=1$, we find $\left\langle{\delta h}\right\rangle\sim r^{0.9}$, which is shallower compared to the $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ scaling for the energy. We have $\left\langle{\delta h}\right\rangle\sim r^{2}$ as $r\rightarrow 0$, similar to the velocity structure function. Overall, the results show that helicity is mainly concentrated at the large scales for all $De$: turbulence cascade tends to restore the mirror symmetry at the small scales in both Newtonian and polymeric turbulence (Ditlevsen & Giuliani, 2001; Chen et al., 2003a; Baj et al., 2022); see §4.3 for additional discussion.
Figure 3: Instantaneous visualisations of the dissipation fields and the polymeric source/sink terms for $De=1$. Top: Dissipation fields $\varepsilon_{f}^{\delta q^{2}}$ (left), $\varepsilon_{f}^{\delta\omega^{2}}$ (centre) and $\varepsilon_{f}^{\delta h}$ (right). Bottom: Polymeric source/sink terms $\pi^{\delta q^{2}}$ (left), $\pi^{\delta\omega^{2}}$ (centre) and $\pi^{\delta h}$ (right). All the quantities are normalised with the modulus of their averaged value.
We now turn our attention to the dissipation rate fields and the polymeric source/sink terms. As illustrated in figure 3 through instantaneous visualisations, the fluid and polymer contributions exhibit markedly different behaviours. For both energy and enstrophy, the fluid dissipation fields $\varepsilon_{f}^{\delta q^{2}}$ and $\varepsilon_{f}^{\delta\omega^{2}}$ are strictly positive, reflecting the fact that viscous effects irreversibly dissipate these quantities at all spatial locations. In contrast, the polymeric terms $\pi^{\delta q^{2}}$ and $\pi^{\delta\omega^{2}}$ can take on both positive and negative values, indicating that polymers may locally act either as sinks or sources of energy and enstrophy. In the case of energy, this behaviour is consistent with the appearance of the same term (but with opposite sign) in the evolution equation for the free energy of the polymers (De Angelis et al., 2005): it represents a bidirectional exchange of energy between the fluid and the polymeric phases. On average, however, $\langle\pi^{\delta q^{2}}\rangle>0$, indicating a net transfer of energy from the fluid to the polymeric phase, where it is eventually dissipated. Figure 3 shows that both positive and negative intense events of $\pi^{\delta\omega^{2}}$ are concentrated in regions of intense shear. Unlike energy and enstrophy dissipation, both $\varepsilon_{f}^{\delta h}$ and $\pi^{\delta h}$ associated with helicity are not sign-definite, meaning they can contribute positively or negatively depending on the local flow configuration.
Refer to captionRefer to caption
@@
We now present a statistical quantification of the fluid dissipation associated with all three quantities in figure 4. These quantities play a key role in the scale-by-scale cascades. In Newtonian HIT, for the two averaged, inviscid invariants (energy and helicity) the dissipation rate matches the injection rate at the large scales and the averaged total flux in the inertial range of scales. The distributions of the dissipations show large tails, which are indicative of small-scale intermittency (Sreenivasan & Antonia, 1997), see figure 3. As expected, the distributions narrow for the polymeric case: polymers reduce the magnitude of the velocity derivatives, thus largely reducing the regions of large dissipations (Liberzon et al., 2005, 2006). The distributions of the energy and enstrophy dissipation collapse nicely for all $De$ once the quantities are normalised with the respective standard deviation, similarly to what was shown by Perlekar et al. (2010) and is consistent with the $De$ invariance of intermittency corrections observed in Rosti et al. (2023).
Refer to captionRefer to caption
@@
Figure 5: Distribution of the the $\pi$ terms for $\delta q^{2}$, $\delta\omega^{2}$ and $\delta h$ for different Deborah numbers. Left: $\pi^{\delta q^{2}}$, centre: $\pi^{\delta\omega^{2}}$ and right: $\pi^{\delta h}$. The bottom panels plot the distribution of the three quantities normalised with their root-mean-square value.
@@
We conclude this section with the distributions of $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$ and $\pi^{\delta h}$. As mentioned above, these quantities are not positive definite, meaning that locally polymers may act as either a sink or a source for the velocity/vorticity fluctuations. The distribution of $\pi^{\delta q^{2}}$ is right-skewed for all $De$, confirming the positive average value $\langle\pi^{\delta q^{2}}\rangle=\langle\epsilon_{p}\rangle>0$. As $De$ increases, both the right and left tails widen, meaning that sink and source intense events are equally promoted by the larger stretching of the polymers. The effect of $De$ is much more significant on $\pi^{\delta\omega^{2}}$, whose distributions are shown in the central panels of figure 5, see also table 1. At small $De<1$, the distributions are right skewed and, on average, the polymers act as a sink for the enstrophy $\langle\pi^{\delta\omega^{2}}\rangle>0$. For $De\geq 1$, instead, the distribution of $\pi^{\delta\omega^{2}}$ is left skewed. In this case, due to the larger stretching of the polymers, the fluid-polymer interaction facilitates intense production of enstrophy $\langle\pi^{\delta\omega^{2}}\rangle<0$ and for large enough $De$ it emerges as the dominant source of enstrophy generation; this will be further discussed in §4.2 and §5. Eventually, the right panels in figure 5 shows that the distribution of $\pi^{\delta h}$ is symmetric, and that both tails become longer as $De$ increases.
4 Scale-by-scale balance equations
4.1 Scale energy $\left\langle{\delta q^{2}}\right\rangle$